# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

### v3

@@ VI Numerical Validation and Simulations
[…]
@@ VII-A Large Language Models
Parameter Settings. Average logical depth: $\mathbb{E}[\text{Dd}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
@@ VII-B Knowledge Graphs
TABLE IV: Query Patterns in the Biomedical KB
@@ VIII-A Information Theory and Kolmogorov Complexity
Shannon’s seminal 1948 paper [shannon1948mathematical] introduced entropy $H(X)=-\sum_{x}p(x)\log p(x)$ as the fundamental measure of information content, establishing the mathematical foundations for data compression and communication. This framework has been extensively developed, with Cover and Thomas [cover2012elements] providing the standard modern treatment encompassing source coding, channel capacity, and rate-distortion theory.
@@
    Physical grounding: We connect logical depth to thermodynamic energy costs via Landauer’s principle, providing physical constraints on computation.
@@ VIII-B Thermodynamics of Computation
    Logical depth-energy connection: We prove that each inference step requires energy $k_{B}T\ln(2)$ by connecting logical depth to bit erasure.
@@ VIII-C Time-Space Trade-offs in Computation
Our work differs by: (i) adopting an information-theoretic perspective quantifying information content rather than worst-case complexity, (ii) incorporating thermodynamic energy costs via Landauer’s principle, (iii) optimizing average-case performance over query distributions using logical depth, and (iv) providing asymptotic bounds rather than competitive ratios. The constant $(1+\ln(2))$ emerges from information-theoretic first principles rather than algorithm-specific analysis.
@@ VIII-E Knowledge Graphs and Semantic Systems
Large-scale knowledge graphs [bollacker2008freebase, carlson2010toward] and semantic networks [miller1995wordnet] face significant storage-computation challenges. Embedding methods [bordes2013translating] compress knowledge at the cost of inexact retrieval. Our framework quantifies this trade-off: compression reduces storage entropy but increases logical depth for reconstruction.
@@ VIII-I Information-Theoretic Learning and Prediction
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with logical depth $\text{Dd}(q|S)$, requiring depth-aware algorithms.
@@ IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental assumption that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Logical Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
@@
This theoretical finding has profound implications for the design of next-generation AI architectures. As current models like Transformers approach the limits of memory bandwidth and energy consumption [vaswani2017attention, horowitz20141], our framework suggests that future systems may benefit from optimizing the Derivation-Retrieval Ratio rather than simply scaling parameters. Such systems could dynamically navigate the ”Logical Depth vs. Shannon Entropy” curve, acting as adaptive engines that minimize the total action of the ontological mapping process.

### v4

@@ VI Numerical Validation and Simulations
Thm 4 (Scaling) Ratio stable at 0.22
Thm 6 (Lower Bound) $\approx$ Valid for large $H(Q|S_{O})$
Polynomial Depth $D_{d}/\log(m)\approx 0.6$ stable
Phase Transition $1.1\%$ theory-empirical gap
FreqDepth Strategy 26% cost reduction vs. LRU
$H(Q)\sim\log(n)$ Scaling Ratio stable at 0.53
The experiments confirm that: (1) the frequency-weighted upper bound (Theorem 5) achieves 100% satisfaction with tightness 0.89 for high-frequency queries; (2) the N-sweep analysis validates the theoretical prediction that optimal storage fraction $\beta^{*}$ increases from 0% to 92% as query horizon $N$ grows from 100 to 50,000; (3) the derivation entropy scaling (Theorem 4) is validated with the ratio $H_{\mathrm{derive}}\cdot L_{S}/H_{q}$ stable at 0.22 across all $m$ values; (4) the information lower bound (Theorem 6) is asymptotically valid but requires sufficiently large $H(Q|S_{O})$ to dominate the $O(1)$ correction term; (5) derivation depth satisfies the polynomial-depth assumption with $D_{d}/\log(m)\approx 0.6$ stable; (6) the marginal cost-benefit analysis predicts the optimal storage fraction $\beta^{*}$ with $1.1\%$ error; and (7) our FreqDepth caching strategy achieves 26% cost reduction over LRU.
@@ VII-A Large Language Models
Parameter Settings. Average derivation depth: $\mathbb{E}[\text{Dd}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
@@ VII-B Knowledge Graphs
TABLE VI: Query Patterns in the Biomedical KB
@@ VIII-A Information Theory and Kolmogorov Complexity
Shannon’s seminal 1948 paper [shannon1948mathematical] introduced entropy $H(X)=-\sum_{x}p(x)\log p(x)$ as the fundamental measure of information content, establishing the mathematical foundations for data compression and communication. This framework has been extensively developed, with Cover and Thomas [cover2006elements] providing the standard modern treatment encompassing source coding, channel capacity, and rate-distortion theory.
@@
    Physical grounding: We connect derivation depth to thermodynamic energy costs via Landauer’s principle, providing physical constraints on computation.
@@ VIII-B Thermodynamics of Computation
    Derivation depth-energy connection: We prove that each inference step requires energy $k_{B}T\ln(2)$ by connecting derivation depth to bit erasure.
@@ VIII-C Time-Space Trade-offs in Computation
Our work differs by: (i) adopting an information-theoretic perspective quantifying information content rather than worst-case complexity, (ii) incorporating thermodynamic energy costs via Landauer’s principle, (iii) optimizing average-case performance over query distributions using derivation depth, and (iv) providing asymptotic bounds rather than competitive ratios. The constant $(1+\ln(2))$ emerges from information-theoretic first principles rather than algorithm-specific analysis.
@@ VIII-E Knowledge Graphs and Semantic Systems
Large-scale knowledge graphs [bollacker2008freebase, carlson2010toward] and semantic networks [miller1995wordnet] face significant storage-computation challenges. Embedding methods [bordes2013translating] compress knowledge at the cost of inexact retrieval. Our framework quantifies this trade-off: compression reduces storage entropy but increases derivation depth for reconstruction.
@@ VIII-I Information-Theoretic Learning and Prediction
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with derivation depth $\text{Dd}(q|S)$, requiring depth-aware algorithms.
@@ IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental assumption that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Derivation Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
@@
This theoretical finding has profound implications for the design of next-generation AI architectures. As current models like Transformers approach the limits of memory bandwidth and energy consumption [vaswani2017attention, horowitz20141], our framework suggests that future systems may benefit from optimizing the Derivation-Retrieval Ratio rather than simply scaling parameters. Such systems could dynamically navigate the ”Derivation Depth vs. Shannon Entropy” curve, acting as adaptive engines that minimize the total action of the ontological mapping process.